7. A company issued 6-year, 8% bonds with a par value of $550,000. The market rate when the bonds were issued was 7.5%. The company received $555,500 cash for the bonds. Using the straight-line method, the amount of recorded interest expense for the first semiannual interest period is:

8. A corporation issued 8% bonds with a par value of $1,010,000, receiving a $22,000 premium. On the interest date 5 years later, after the bond interest was paid and after 40% of the premium had been amortized, the corporation purchased the entire issue on the open market at 99 and retired it. The gain or loss on this retirement is:

20.Prior to May 1, Fortune Company has never had any treasury stock transactions. A company repurchased 220 shares of its common stock on May 1 for $11,000. On July 1, it reissued 110 of these shares at $52 per share. On August 1, it reissued the remaining treasury shares at $49 per share. What is the balance in the Paid-in Capital, Treasury Stock account on August 2?

21.A company had net cash flows from operations of $142,000, cash flows from financing of $374,000, total cash flows of $566,000, and average total assets of $3,820,000. The cash flow on total assets ratio equals

23. A company's income statement showed the following: net income, $139,000 and depreciation expense, $34,500. An examination of the company's current assets and current liabilities showed the following changes as a result of operating activities: accounts receivable decreased $10,900; merchandise inventory increased $21,000; and accounts payable increased $4,900. Calculate the net cash provided or used by operating activities.
Answer #1

7) Interest expense = (550000*4%)-(555500-550000/12) = $21542

8) Gain or loss on retirement = (1010000+22000*60%)-(1010000*90%) = 114200

Loss on retirement = $114200

20) Paid in capital from treasury stock = (110*2)-(110*1) = $110

21) Cash flow to total assets = 142000/3820000 = 0.04 Times

23) Net cash flow from operating activities = 139000+34500+10900-21000+4900 = $168300